# Variational Calculation of the Hyperfine Stark Effect in Atomic 87Rb, 133Cs, and 169Tm

### v1

@@ Abstract
An electronically variational approach to the calculation of atomic hyperfine structure transition energies under the influence of static external electric fields is presented. The method avoids the calculation of intermediate atomic states entirely and requires only the wavefunctions of the electronic states involved in the respective hyperfine levels. These wavefunctions are obtained through relativistic general-excitation-rank configuration interaction theory. The method also enables for calculations on atoms with the most complicated of shell structures.
The initial applications include 87Rb and 133Cs where very good agreement of the approach with literature results is established. For 169Tm that is used in the development of atomic clocks the differential static electric dipole polarizability between ground levels $J=\frac{7}{2}$ and $J=\frac{5}{2}$ is calculated to be $\Delta\alpha=-0.23\pm 0.11$ a.u. The hyperfine Stark coefficient for the hyperfine levels belonging to the ground term with $J=\frac{7}{2}$ is found to be $k=(1.3\pm 1.0)\times 10^{-13}$ [Hz/((V/m)2)]. This coefficient is several orders of magnitude smaller than the corresponding coefficients in 87Rb and 133Cs.
@@ I Introduction
Indeed, hyperfine-level transitions in the thulium atom have recently been used as clock transitions in the design of an optical atomic clock with unusually low sensitivity to the BBR NatComm_Tm_2019 ; PRA_Tm_2016 ; PRA_Tm_2020 . The relevant BBR frequency shift has been experimentally estimated NatComm_Tm_2019 to be a few orders of magnitude smaller than the corresponding shift in clock transitions of other neutral atoms. The uncertainty of this measurement has been given as around $50$% AIP_Tm_2020 for the electric component of the BBR. An extensive theoretical study of the thulium atom with relevance to its use as an atomic clock has been presented very recently fritzsche_Tm_2024 , but the BBR or hyperfine Stark shifts have not been addressed in that work.
The paper is organized as follows. In section II the theoretical approach is laid out in detail and a comparison with the common perturbative approaches to the calculation of hyperfine Stark shifts is drawn. In the following section III the hyperfine Stark coefficients for three atoms of relevance to atomic-clock research are presented using the present method. In addition, the differential static dipole polarizability relevant to the clock transition in 169Tm is calculated using a finite-field approach that has earlier been applied to the calculation of spin-orbit resolved electric dipole polarizabilities in atomic states fleig_halo ; fleig_pol13 . In the final section IV conclusions from the presented work are drawn.
@@ II.1 Static Hyperfine Stark Shift
$\Delta\varepsilon(E_{\text{ext}})=\varepsilon_{J_{u},{M_{J}}_{u}}(E_{\text{ext}})+\varepsilon_{F_{u}}(E_{\text{ext}})-\big{(}\varepsilon_{J_{l},{M_{J}}_{l}}(E_{\text{ext}})+\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}.$ (1)
where $J$ denotes total electronic angular momentum associated with the respective hyperfine level. $\varepsilon_{F}$ is thus the hyperfine energy relative to the respective electronic reference energy $\varepsilon_{J,M_{J}}$. Note that generally $J_{u}\neq J_{l}$.
@@
$k=\frac{\delta\nu(E_{\text{ext}})}{E_{\text{ext}}^{2}}$ $=$ $\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{J_{u},{M_{J}}_{u}}(E_{\text{ext}})+\varepsilon_{F_{u}}(E_{\text{ext}})-\varepsilon_{J_{l},{M_{J}}_{l}}(E_{\text{ext}})-\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}\right.$ (5)
$\hskip 42.67912pt\left.-\big{(}\varepsilon_{J_{u},{M_{J}}_{u}}(0)+\varepsilon_{F_{u}}(0)-\varepsilon_{J_{l},{M_{J}}_{l}}(0)-\varepsilon_{F_{l}}(0)\big{)}\right]$
$=$ $\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{J_{u},{M_{J}}_{u}}(E_{\text{ext}})-\varepsilon_{J_{l},{M_{J}}_{l}}(E_{\text{ext}})\big{)}+\big{(}\varepsilon_{F_{u}}(E_{\text{ext}})-\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}\right.$
$\hskip 42.67912pt\left.-\big{(}\varepsilon_{J_{u},{M_{J}}_{u}}(0)-\varepsilon_{J_{l},{M_{J}}_{l}}(0)\big{)}-\big{(}\varepsilon_{F_{u}}(0)-\varepsilon_{F_{l}}(0)\big{)}\right]$
@@
    $J_{u}\neq J_{l}$, i.e., the two respective hyperfine levels belong to different electronic states. Then the Stark coefficient is stongly dominated by the electric polarizability difference between the different electronic wavefunctions and it can be written as
    $k\approx\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{J_{u},{M_{J}}_{u}}(E_{\text{ext}})-\varepsilon_{J_{l},{M_{J}}_{l}}(E_{\text{ext}})\big{)}-\big{(}\varepsilon_{J_{u},{M_{J}}_{u}}(0)-\varepsilon_{J_{l},{M_{J}}_{l}}(0)\big{)}\right]$ (6)
@@
    $J_{u}=J_{l}$, i.e., the two respective hyperfine levels belong to the same electronic states. In this case all of the electronic energies in the expression (5) cancel pairwise and the Stark coefficient is
@@ II.2 Field-Dependent Hyperfine Level Energy
$\hat{H}({{E}_{\text{ext}}})$ ::=:= $\hat{H}^{\text{Dirac-Coulomb}}+\hat{H}^{\text{Int-Dipole}}$ (12)
@@ II.3 Comparison with Perturbation Theory
where $V_{kl}=\left<\psi_{k}^{(0)}|\hat{V}|\psi_{l}^{(0)}\right>$ and $\varepsilon_{m}$ is an unperturbed energy. If terms ${\cal{O}}(\hat{V}^{3})$ are omitted from Eq. (15) and the resulting truncated expansion is used to evaluate the expectation value $\psi_{0}\right|\hat{H}_{\text{HF}}\left|\psi_{0}\right>$ then the third-order expressions from Refs. schwartz_SPT_1959 and Angstmann_Dzuba_Flambaum_PRA2006 arise. In the present work, however, ${\cal{O}}(\hat{V}^{3})$ is implicitly included in Eq. (11) which leads to the presence of all orders of the external electric field in Eq. (9) and, therefore, also in Eq. (5) defining the hyperfine Stark coefficient.
@@ III.1.1 External Electric Field
The external electric field $E_{z}$ has to be chosen small enough to assure that $k$ as defined in Eq. (5) does not vary with $E_{z}$, but also large enough to ensure that the field-dependent change in $\psi(E_{z})$ is sufficiently greater than the convergence threshold chosen for the wavefunction. Typically, this is achieved for $E_{z}=10^{-3}$ a.u.  which is the value used in the present applications. In addition, a region around $E_{z}=10^{-3}$ a.u.  has been explored which allows for an estimate of the uncertainty in $k$ due to deviations from quadratic dependency of the field-dependent total energy.
@@ III.1.2 Atomic Basis Sets
For all three atoms Gaussian basis sets are used. The Rb atom is described by the uncontracted aug-cc-pwCVQZ-X2C set from the EMSL library EMSL-basis2019 that has been singly densified in the $s$ and $p$ spaces and augmented by one dense and one diffuse function as described in Ref. Hubert_Fleig_2022 . The full uncontracted set then comprises $77s,59p,21d,5f,3g$ functions. The Cs atom is described by the uncontracted Dyall vQZ basis dyall_s-basis with $5s5p6s6p$ correlating exponents added. Like for Rb the $s$ and $p$ spaces have been densified once and augmented. The final uncontracted set comprises $75s,61p,19d,3f,1g$ functions.
For the thulium atom three different basis sets dyall_4f are employed: 1) The uncontracted Dyall vTZ set including all $[4f/6s/5s,p,d]$ correlating and $4f$ dipole-polarizing functions adding up to $[30s,24p,18d,13f,4g,2h]$ functions (in the following denoted as TZ). 2) The uncontracted Dyall cvQZ basis including all $[4f/6s/5d]$ correlating and $4f$ dipole-polarizing functions comprised by a total of $[35s,30p,19d,16f,6g,4h,2i]$ functions (QZ) and 3) The latter basis set with the $s$ and $p$ spaces simply densified and augmented as described in Ref. Hubert_Fleig_2022 , yielding $[71s,61p,19d,16f,6g,4h,2i]$ primitive functions (QZ+). The densification and augmentation in the $s$ and $p$ spaces serves to produce a highly accurate basis set for describing the $s$-$p$ mixing which is predominantly responsible for the energy shifts when an external electric field is applied to hyperfine states.
@@ III.1.3 Atomic Wavefunctions
All atoms considered here have an odd number of electrons. In addition, the interaction with $E_{z}$ breaks the full rotational symmetry of the atom which means that only $M_{J}$ remains as an exact electronic quantum number (for $\psi(E_{\text{ext}})$ in Eq. (11)). Atomic wavefunctions are calculated in the $\left|M_{J}\right|=\Omega=\frac{1}{2}$ irreducible representation of the $C_{\infty v}^{*}$ double point group. The target atom is placed at the origin of the reference frame and a ghost atom with neither electronic basis functions nor nuclear charge is placed at a finite distance along the axis of the external $E$ field in order to allow for the inclusion of the external-field Hamiltonian.
@@
Acronyms are used for brevity in defining atomic correlated wavefunctions from the second step. As an example, SDT9_10au stands for Single, Double, Triple replacements relative to the DCHF reference state where 9 electrons occupying the outermost shells in the DCHF reference state are taken into account in the correlation expansion and the complementary space of virtual spinors is truncated at 10 a.u. An acronym SD8_SDT9 means that up to two holes are allowed in the shells occupied by 8 electrons and up to triple excitations into the virtual spinors are allowed from the combined shells occupied by the 9 electrons in the reference state.

### v2

@@ Abstract
An electronically variational approach to the calculation of atomic hyperfine structure transition energies under the influence of static external electric fields is presented. The method avoids the calculation of intermediate atomic states entirely and requires only the wavefunctions of the electronic states involved in the respective hyperfine levels. These wavefunctions are obtained through relativistic general-excitation-rank configuration interaction theory. A variant of the method also enables for calculations on atoms with the most complicated of shell structures.
Applications to 87Rb, 133Cs and a specific clock transition in 169Tm are presented. The final results $k_{\text{Rb}}=-1.234\pm 0.0223$ $10^{-10}$ Hz/((V/m)2)] and $k_{\text{Cs}}=-2.347\pm 0.084$ $10^{-10}$ Hz/((V/m)2)] obtained under inclusion of up to quintuple excitations in the atomic wavefunction expansion are compatible with previous calculations and, in the case of Cs, confirm that one of the earlier experimental measurements is not reliable. For 169Tm that is used in the development of atomic clocks the differential static scalar electric dipole polarizability between ground levels $J=\frac{7}{2}$ and $J=\frac{5}{2}$ is calculated to be $\Delta\alpha^{s}_{0}=-0.134\pm 0.11$ a.u. This result from a pure ab initio calculation confirms the result of $\Delta\alpha^{s}_{0}=-0.063^{+0.01}_{-0.005}$ a.u. obtained in Nat. Comm. 10 (2019) 1724 where a combination of measurement and theoretical modeling has been used.
@@ I Introduction
Indeed, hyperfine-level transitions in the thulium atom have recently been used as clock transitions in the design of an optical atomic clock with unusually low sensitivity to the BBR NatComm_Tm_2019 ; PRA_Tm_2016 ; PRA_Tm_2020 . The relevant BBR frequency shift has been experimentally estimated NatComm_Tm_2019 to be a few orders of magnitude smaller than the corresponding shift in clock transitions of other neutral atoms. The uncertainty of this measurement has been given as around 5050% AIP_Tm_2020 for the electric component of the BBR. An extensive theoretical study of the thulium atom with relevance to its use as an atomic clock has been presented very recently fritzsche_Tm_2024 , but the BBR or hyperfine Stark shifts have not been addressed in that work.
The paper is organized as follows. In section II the theoretical approach is laid out in detail and a comparison with the common perturbative approaches to the calculation of hyperfine Stark shifts is drawn. In the following section III the hyperfine Stark coefficients for three atoms of relevance to atomic-clock research are presented using the present methods. In addition, the differential static dipole polarizability relevant to the clock transition in 169Tm is calculated using a finite-field approach that has earlier been applied to the calculation of spin-orbit resolved electric dipole polarizabilities in atomic states fleig_halo ; fleig_pol13 . In the final section IV conclusions from the presented work are drawn.
@@ II Theory
The present approach differs substantially from previous approaches Angstmann_Dzuba_Flambaum_PRA2006 ; Safronova_BBR_2009 ; beloy_derevianko_PRL2006 to calculate electric polarizabilities of atoms including the hyperfine interaction. Before presenting the formal details it is, therefore, in place to lay out the big picture of the present method.
In a hypothetical atom without relativistic effects – in particular without the spin-orbit interaction – polarizabilities can be defined for orbital angular momentum ($L$) microstates and denoted $\alpha_{L,M_{L}}$. If now very weak spin-orbit interaction is taken into account then the total angular momentum $J$ becomes an exact quantum number. Under the assumption, however, that the magnetic coupling is very weak, $M_{L}$ remains an approximately valid quantum number and the corresponding polarizabilities can be labelled $\alpha_{J,L,(M_{L})}$. For clarity of notation, the approximate quantum number is set in parentheses.
By analogy, polarizabilities for atoms without magnetic hyperfine interaction but including the electronic spin-orbit interaction can be defined exactly as $\alpha_{J,M_{J}}$. Including now a very weak magnetic hyperfine interaction leaves $M_{J}$ as an approximately valid quantum number and the polarizabilities are correspondingly labelled $\alpha_{F,J,(M_{J})}$ where $F\in\{J+I,\ldots,|J-I|\}$ and $I$ is the nuclear spin quantum number. This is the picture adopted in the present approach.
In terms of methodology, the first step consists in variationally determining the wavefunctions of the atomic electronic states in question including relativistic effects, electron correlation effects, and fully relaxing the wavefunction with respect to the externally applied finite uniform electric field. In the second step, the obtained wavefunctions are used as zeroth-order wavefunctions in a perturbative first-order evaluation of the hyperfine interaction in the relevant atomic states.
@@ II.1 Static Hyperfine Stark Shift
The following formulation is energy-based and in its present form only applicable to atomic $S$ (orbital angular momentum free) states where non-scalar static electric polarizability is zero. The more general case will be presented in forthcoming work. Thus, the present formulation is applied to Rubidium and Cesium electronic ground states. For Thulium, a different approach is chosen as to be discussed below.
@@
$\Delta\varepsilon(E_{\text{ext}})=\varepsilon_{J_{u},({M_{J}}_{u})}(E_{\text{ext}})+\varepsilon_{F_{u}}(E_{\text{ext}})-\big{(}\varepsilon_{J_{l},({M_{J}}_{l})}(E_{\text{ext}})+\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}.$ (1)
where $J$ denotes total electronic angular momentum associated with the respective hyperfine level. $\varepsilon_{F}$ is thus the hyperfine energy relative to the respective electronic reference energy $\varepsilon_{J,M_{J}}$.
@@
$k=\frac{\delta\nu(E_{\text{ext}})}{E_{\text{ext}}^{2}}$ $=$ $\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{J_{u},({M_{J}}_{u})}(E_{\text{ext}})+\varepsilon_{F_{u}}(E_{\text{ext}})-\varepsilon_{J_{l},({M_{J}}_{l})}(E_{\text{ext}})-\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}\right$. (5)
$\hskip 42.67912pt\left.-\big{(}\varepsilon_{J_{u},({M_{J}}_{u})}(0)+\varepsilon_{F_{u}}(0)-\varepsilon_{J_{l},({M_{J}}_{l})}(0)-\varepsilon_{F_{l}}(0)\big{)}\right]$
$=$ $\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{J_{u},({M_{J}}_{u})}(E_{\text{ext}})-\varepsilon_{J_{l},({M_{J}}_{l})}(E_{\text{ext}})\big{)}+\big{(}\varepsilon_{F_{u}}(E_{\text{ext}})-\varepsilon_{F_{l}}(E_{\text{ext}})\big{)}\right$.
$\hskip 42.67912pt\left.-\big{(}\varepsilon_{J_{u},({M_{J}}_{u})}(0)-\varepsilon_{J_{l},({M_{J}}_{l})}(0)\big{)}-\big{(}\varepsilon_{F_{u}}(0)-\varepsilon_{F_{l}}(0)\big{)}\right]$
@@
    The two respective hyperfine levels belong to different electronic states. Then the Stark coefficient is stongly dominated by the electric polarizability difference between the different electronic wavefunctions and it can be written as
    $k\approx\frac{1}{hE_{\text{ext}}^{2}}\left[\big{(}\varepsilon_{J_{u},({M_{J}}_{u})}(E_{\text{ext}})-\varepsilon_{J_{l},({M_{J}}_{l})}(E_{\text{ext}})\big{)}-\big{(}\varepsilon_{J_{u},({M_{J}}_{u})}(0)-\varepsilon_{J_{l},({M_{J}}_{l})}(0)\big{)}\right]$ (6)
@@
    The two respective hyperfine levels belong to the same electronic state. In this case all of the electronic energies in the expression (5) cancel pairwise and the Stark coefficient is
@@ II.2 Field-Dependent Hyperfine Level Energy
$\hat{H}({{E}_{\text{ext}}})$ $:=$ $\hat{H}^{\text{Dirac-Coulomb}}+\hat{H}^{\text{Int-Dipole}}$ (12)
@@ II.3 Comparison with Perturbation Theory
where $V_{kl}=\left<\psi_{k}^{(0)}|\hat{V}|\psi_{l}^{(0)}\right$> and $\varepsilon_{m}$ is an unperturbed energy. If terms ${\cal{O}}(\hat{V}^{3})$ are omitted from Eq. (15) and the resulting truncated expansion is used to evaluate the expectation value $\left<\psi_{0}\right|\hat{H}_{\text{HF}}\left|\psi_{0}\right$> then the third-order expressions from Refs. Angstmann_Dzuba_Flambaum_PRA2006 and schwartz_SPT_1959 arise. In the present work, however, ${\cal{O}}(\hat{V}^{3})$ is implicitly included in Eq. (11) which leads to the presence of all orders of the external electric field in Eq. (9) and, therefore, also in Eq. (5) defining the hyperfine Stark coefficient.
II.4 Scalar and Tensor Static Polarizabilites
An electric dipole polarizability for a state labelled $J,M_{J}$ can be given angel_sandars_pol1 ; Manakov_Ovsyannikov_1979 ; Chen_Raithel_2015 in terms of scalar ($\alpha_{0}$) and tensor ($\alpha_{2}$) static polarizabilites as follows:
$\alpha_{J,M_{J}}=\alpha_{0}(J)+\alpha_{2}(J)\frac{3M_{J}^{2}-J(J+1)}{J(2J-1)}$ (16)
where $\alpha_{0}$ and $\alpha_{2}$ are functions of $J$ only. If $\alpha_{J,M_{J}}$ and $\alpha_{J,M_{J}^{\prime}}$ with $M_{J}\neq M_{J}^{\prime}$ are known then the resulting system of Eqs. (16) can be inverted and $\alpha_{0}$ and $\alpha_{2}$ become calculable functions of $\alpha_{J,M_{J}}$ and $\alpha_{J,M_{J}^{\prime}}$. This fact will be exploited in order to compare with polarizability results given in terms of scalar and tensor polarizabilities in cases where $J>\frac{1}{2}$.
This is the case for the Thulium atom ground levels. In the calculations presented below $\alpha_{J,M_{J}}$ are calculated using the finite-field method (see subsection III.1.4). $\alpha_{0}(J)$ and $\alpha_{2}(J)$ are then obtained by inverting Eqs. (16). Finally, the differential static scalar polarizability in this case results from
$\Delta\alpha_{0}=\alpha_{0}(J_{u})-\alpha_{0}(J_{l})$ (17)
and the hyperfine Stark shift parameter can be obtained from this as
$k=-\frac{1}{2}\,\Delta\alpha_{0}$ (18)
@@ III.1.1 External Electric Field
The external electric field $E_{z}$ has to be chosen small enough to assure that $k$ as defined in Eq. (5) does not vary with $E_{z}$, but also large enough to ensure that the field-dependent change in $\psi(E_{z})$ is sufficiently greater than the convergence threshold chosen for the wavefunction. Typically, this is achieved for $E_{z}=10^{-3}$ a.u.  which is the value used in the present direct calculations. In addition, a region around $E_{z}=10^{-3}$ a.u.  has been explored which allows for an estimate of the uncertainty in $k$ due to deviations from quadratic dependency of the field-dependent total energy.
@@ III.1.2 Atomic Basis Sets
For all three atoms Gaussian basis sets are used. The Rb atom is described by the uncontracted Dyall vTZ set with $28s,20p,12d,1f$ primitive functions and the aug-cc-pwCVQZ-X2C set from the EMSL library EMSL-basis2019 that has been singly densified in the $s$ and $p$ spaces and augmented by one dense and one diffuse function as described in Ref. Hubert_Fleig_2022 . The full uncontracted QZ+ set then comprises 77s,59p,21d,5f,3g77s,59p,21d,5f,3g functions. The Cs atom is described by the uncontracted Dyall vTZ set with $31s,24p,15d,1f$ primitive functions and the vQZ basis dyall_s-basis with $5s5p6s6p$ correlating exponents added. Like for Rb the $s$ and $p$ spaces have been densified once and augmented. The final uncontracted QZ+ set comprises $75s,61p,19d,3f,1g$ functions.
For the thulium atom two different basis sets dyall_4f are employed: 1) The uncontracted Dyall vTZ set including all $[4f/6s/5s,p,d]$ correlating and $4f$ dipole-polarizing functions adding up to $[30s,24p,18d,13f,4g,2h]$ functions (in the following denoted as TZ). 2) The uncontracted Dyall cvQZ basis including all $[4f/6s/5d]$ correlating and $4f$ dipole-polarizing functions comprised by a total of $[35s,30p,19d,16f,6g,4h,2i]$ functions (QZ).
The densification and augmentation in the $s$ and $p$ spaces in some cases serves to test a highly accurate basis set for describing the $s$-$p$ mixing which is predominantly responsible for the energy shifts when an external electric field is applied to hyperfine states.
@@ III.1.3 Atomic Wavefunctions
All atoms considered here have an odd number of electrons. In addition, the interaction with $E_{z}$ breaks the full rotational symmetry of the atom which means that only $M_{J}$ remains as an exact electronic quantum number (for $\psi(E_{\text{ext}})$ in Eq. (11)). Atomic wavefunctions are calculated in the $\left|M_{J}\right|=\Omega$ irreducible representations of the $C_{\infty v}^{*}$ double point group. The target atom is placed at the origin of the reference frame and a ghost atom with neither electronic basis functions nor nuclear charge is placed at a finite distance along the axis of the external $E$ field in order to allow for the inclusion of the external-field Hamiltonian in linear-symmetry calculations. The ghost atom introduces no physical interaction.
@@
Acronyms are used for brevity in defining atomic correlated wavefunctions from the second step. As an example, SDT9_10au stands for Single, Double, Triple replacements relative to the DCHF reference state where 9 electrons occupying the outermost shells in the DCHF reference state are taken into account in the correlation expansion and the complementary space of virtual spinors is truncated at 10 a.u. An acronym SD8_SDT9 means that up to two holes are allowed in the shells occupied by 8 electrons and up to triple excitations into the virtual spinors are allowed from the combined shells occupied by the 9 electrons in the reference state. In the case of an alkali atom this means that up to double excitations from the (n1)s(n1)p(n-1)s~(n-1)p shells and up to triple excitations from the combined (n1)s(n1)p(n-1)s~(n-1)p and $ns$ shells are included in the wavefunction expansion. Thus, the model SD8_SDT9 comprises a subset of the determinants of the model SDT9, where in the latter all triple excitations from the (n1)s(n1)p(n-1)s~(n-1)p and $ns$ shells are included.
@@ III.1.4 Finite-Field (FF) Method
III.2.1 Static Electric Dipole Polarizability $\alpha_{D}$
As a corroboration of the present method of calculating polarizabilites a comparison with literature results for scalar and tensor polarizabilities is drawn. Using Eq. (16) the $M_{J}$-dependent values of $\alpha_{D}$ are obtained from the equations
$\alpha_{D}\left({}^{2}P_{3/2,1/2}\right)$ $=$ $\alpha_{0}\left({}^{2}P_{3/2}\right)-\alpha_{2}\left({}^{2}P_{3/2}\right)$
$\alpha_{D}\left({}^{2}P_{3/2,3/2}\right)$ $=$ $\alpha_{0}\left({}^{2}P_{3/2}\right)+\alpha_{2}\left({}^{2}P_{3/2}\right)$ (19)
where $\alpha_{0}$ is the scalar polarizability and $\alpha_{2}$ is the tensor polarizability. This allows for calculating the $M_{J}$-dependent polarizabilities from the results in Ref. Arora_Sahoo_2012 and these are given in Table 1. The inverted equations (19) read
$\alpha_{0}\left({}^{2}P_{3/2}\right)$ $=$ $\frac{1}{2}\left[\alpha_{D}\left({}^{2}P_{3/2,3/2}\right)+\alpha_{D}\left({}^{2}P_{3/2,1/2}\right)\right]$
$\alpha_{2}\left({}^{2}P_{3/2}\right)$ $=$ $\frac{1}{2}\left[\alpha_{D}\left({}^{2}P_{3/2,3/2}\right)-\alpha_{D}\left({}^{2}P_{3/2,1/2}\right)\right]$ (20)
This, in turn, allows for calculating the scalar and tensor polarizabilities from the present $M_{J}$-dependent values, the results of which are also given in Table 1.
Table 1: $M_{J}$-dependent static electric dipole polarizabilities $\alpha_{D}$ [a.u.] calculated through the FF method for states $^{M}L_{J,M_{J}}$ where $M=2S+1$ is the spin multiplicity and scalar ($\alpha_{0}$) and tensor ($\alpha_{2}$) polarizabilities for the $^{2}P_{3/2}$ state; for the electronic ground state $^{2}S_{1/2,1/2}$, $\alpha_{D}\left({}^{2}S_{1/2,1/2}\right)=\alpha_{0}\left({}^{2}S_{1/2}\right)$.
$5s^{1}$ $5p^{1}$
Model $^{2}S_{1/2,1/2}$ $^{2}P_{1/2,1/2}$ $^{2}P_{3/2,1/2}$ $^{2}P_{3/2,3/2}$ $\alpha_{0}(^{2}P_{3/2})$ $\alpha_{2}(^{2}P_{3/2})$
DCHF $485.3$
QZ+/SDT9/10au $333.0$ $820.3$ $1047.9$ $719.3$ $883.6$ 164.3-164.3
Experiment (cited in Ref. Arora_Sahoo_2012 ) $318.79$ $810.6$ $1020$ $694$ $857.0$ 163-163
Other theory (Arora_Sahoo_2012 ) $318.3$ $810.5$ $1033.9$ $702.1$ $868.0$ 165.9-165.9
For the $^{2}P_{3/2}$ level the present results for $\alpha_{0}$ and $\alpha_{2}$ which are derived from $M_{J}$-dependent calculations as described above differ from the experimental values by only 131-3%. Further improvements could be made by modifying the employed electronic-structure model, but for the present case the obtained correspondence serves as a sufficient proof of principle for the applied method.
III.2.2 Hyperfine Stark Effect